# Fractal counterparts of density combinatorics theorems in continued fractions

### v1

@@ Appendix B Arithmetic progressions in ${\rm PS}(\alpha)$
Let $21<\alpha<$ 2. Let $A\subset{\rm PS}(\alpha)$ satisfy $\overline{d}(A|{\rm PS}(\alpha))>0$. There exists a strictly increasing sequence $(N_{k})_{k=1}^{\infty}$ of positive integers such that for every $k\geq 1$,
#$\frac{\#(A\cap[1,N_{k}])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}\geq\frac{\overline{d}(A|{\rm PS}(\alpha))}{2}.$
@@
#$\frac{\#(B\cap[1,\#({\rm PS}(\alpha)\cap[1,N_{k}])])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}=\frac{\#(A\cap[1,N_{k}])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}\geq\frac{\overline{d}(A|{\rm PS}(\alpha))}{2}>0.$
This implies $d_{\rm B}(B)\geq\overline{d}(B)>0$. By [32, Corollary 5], $\{(n,\lfloor n^{\alpha}\rfloor)\in\mathbb{N}^{2}\colon n\in B\}$ contains an $\ell$-term arithmetic progression in $\mathbb{N}^{2}$ for every $\ell\geq 3$. Therefore, $A=\{\lfloor n^{\alpha}\rfloor\colon n\in B\}$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.∎
Appendix C Generalizations to Iterated Function Systems
A countable collection of self-maps is called an Iterated Function System (IFS). It is well-known that the regular continued fraction (1.1) is generated by an IFS on $[0,1]$. Theorems 1.1 and 1.4 can be generalized to a wide class of expansions of real numbers in $[0,1]$ that are generated by IFSs.
Let $\{f_{k}\colon k\in\mathbb{N}\}$ be a collection of $C^{1}$ maps from $[0,1]$ to itself and let $d>1$. We say $\{f_{k}\colon k\in\mathbb{N}\}$ is a $d$-decaying IFS [18] if the following conditions hold:
  • (i)
    there exists $m\in\mathbb{N}$ and $\gamma\in(0,1)$ such that for every $(a_{1},\ldots,a_{m})\in\mathbb{N}^{m}$ and every $x\in[0,1]$ we have $0<|(f_{a_{1}}\circ\cdots\circ f_{a_{m}})^{\prime}(x)|\leq\gamma;$
  • (ii)
    for all distinct $i,j\in\mathbb{N},$ $f_{i}((0,1))\cap f_{j}((0,1))=\emptyset;$
  • (iii)
    for any $\varepsilon>0,$ there exist $c_{1}(\varepsilon)$, $c_{2}(\varepsilon)$ with $0<c_{1}(\varepsilon)\leq 1\leq c_{2}(\varepsilon)$ such that for every $n\in\mathbb{N}$ and every $x\in[0,1]$ we have
    $\frac{c_{1}(\varepsilon)}{k^{d(1+\varepsilon)}}\leq|f_{k}^{\prime}(x)|\leq\frac{c_{2}(\varepsilon)}{k^{d(1-\varepsilon)}};$
  • (iv)
    $\overline{\bigcup_{k=1}^{\infty}f_{k}([0,1])}=[0,1]$, and if $i<j$ then $f_{i}(x)>f_{j}(x)$ for all $x\in[0,1].$
If $\{f_{k}\colon k\in\mathbb{N}\}$ is a $d$-decaying IFS, then by (i) one can define a natural projection $\Pi\colon\mathbb{N}^{\mathbb{N}}\rightarrow[0,1]$ by
$\Pi((a_{n})_{n=1}^{\infty})=\lim_{n\to\infty}f_{a_{1}}\circ\cdots\circ f_{a_{n}}(1).$
For any $x\in[0,1],$ there exists a sequence $(a_{n})_{n=1}^{\infty}$ of positive integers such that
$x=\Pi((a_{n})_{n=1}^{\infty}).$ (C.1)
By (ii), this correspondence is one-to-one up to a countable number of points, and (C.1) gives an expansion of $x$ into an infinite series. The sequence $(a_{n})_{n=1}^{\infty}$ in (C.1) may be viewed as a digit sequence of $x$.
The regular continued fraction is generated by the $2$-decaying IFS $\{f_{k}\colon k\in\mathbb{N}\}$ defined by $f_{k}(x)=1/(x+k)$ with $m=2$. Our setup and main results can be generalized to number expansions generated by arbitrary $d$-decaying IFSs. The constant $1/2$ in Theorem 1.1 should be replaced by $1/d$, and the constant $1/(2\alpha)$ in Theorem 1.4 should be replaced by $1/(d\alpha)$. Lemma 2.1 is the only place where the specific formulas for the regular continued fraction were used. In this generalization, (iii) can be used as a substitute for Lemma 2.1.
Let us stick to Billingsley’s notation for continued fractions. Each $x\in(0,1)\setminus\mathbb{Q}$ has the unique continued fraction expansion of the form
[…]
where $a_{n}\in\mathbb{N}$ for $n\in\mathbb{N}$, which is called the minus or backward continued fraction [16]. This continued fraction is generated by the IFS $\{f_{k}\colon k\in\mathbb{N}\}$ on $[0,1]$ defined by $f_{k}(x)=1-1/(x+k)$ that does not satisfy (i) since $f_{1}(0)=0$ and $f_{1}^{\prime}(0)=1$. Our main results can be generalized to this expansion and the proofs remain intact, since only those $x$ with $a_{n}\to\infty$ as $n\to\infty$ is concerned in the setup. Even further, our main results can be generalized to semi-regular continued fractions that include the regular and the backward continued ones as special cases, see [26] for details.
@@ Acknowledgments
We thank Yasushi Nagai and Kota Saito for their careful readings of an earlier version of the draft and giving valuable suggestions. We thank Nima Alibabaei, Atsushi Katsuda, Shunsuke Usuki for fruitful discussions. This research was supported by the JSPS KAKENHI 23K20220.

### v2

@@ Appendix B Arithmetic progressions in ${\rm PS}(\alpha)$
Let $1<\alpha<2$. Let $A\subset{\rm PS}(\alpha)$ satisfy $\overline{d}(A|{\rm PS}(\alpha))>0$. There exists a strictly increasing sequence $(N_{k})_{k=1}^{\infty}$ of positive integers such that for every $k\geq 1$,
$\frac{\#(A\cap[1,N_{k}])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}\geq\frac{\overline{d}(A|{\rm PS}(\alpha))}{2}.$
@@
$\frac{\#(B\cap[1,\#({\rm PS}(\alpha)\cap[1,N_{k}])])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}=\frac{\#(A\cap[1,N_{k}])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}\geq\frac{\overline{d}(A|{\rm PS}(\alpha))}{2}>0.$
This implies $d_{\rm B}(B)\geq\overline{d}(B)>0$. By [37, Corollary 5], $\{(n,\lfloor n^{\alpha}\rfloor)\in\mathbb{N}^{2}\colon n\in B\}$ contains an $\ell$-term arithmetic progression in $\mathbb{N}^{2}$ for every $\ell\geq 3$, that is, for every $\ell\geq 3$ there exist $n_{1},\ldots,n_{\ell}\in\mathbb{N}$ and $m_{1},m_{2}\in\mathbb{N}$ such that $n_{k}=n_{1}+m_{1}(k-1)$ and $\lfloor n_{k}^{\alpha}\rfloor=\lfloor n_{1}^{\alpha}\rfloor+m_{2}(k-1)$ for $k=2,\ldots,\ell$. Therefore, $A=\{\lfloor n^{\alpha}\rfloor\colon n\in B\}$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.∎
@@ Acknowledgments
We thank anonymous referees for their careful readings of the manuscript and giving valuable comments and suggestions. We thank Yasushi Nagai and Kota Saito for their careful readings of an earlier version of the draft. We thank Nima Alibabaei, Atsushi Katsuda, Shunsuke Usuki for fruitful discussions. YN was supported by the JSPS KAKENHI 25K17282, Grant-in-Aid for Early-Career Scientists. HT was supported by the JSPS KAKENHI 25K21999, Grant-in-Aid for Challenging Research (Exploratory).